# No Other Representation Component Is Needed: Diffusion Transformers Can Provide Representation Guidance by Themselves

### v1

@@ B.2 Stochastic Interpolants Models (SiT)
${\mathbf{x}}_{t}=\alpha_{t}{\mathbf{x}}_{0}+\sigma_{t}{\mathbf{\epsilon}},\quad\alpha_{0}=\sigma_{1}=1,\,\,\alpha_{1}=\sigma_{0}=0,$ (11)
@@
The velocity ${\mathbf{v}}({\mathbf{x}},t)$ is expressed as:
${\mathbf{v}}({\mathbf{x}},t)=\mathbb{E}[\dot{{\mathbf{x}}}_{t}|{\mathbf{x}}_{t}={\mathbf{x}}]=\dot{\alpha}_{t}\mathbb{E}[{\mathbf{x}}_{\ast}|{\mathbf{x}}_{t}={\mathbf{x}}]+\dot{\sigma}_{t}\mathbb{E}[{\mathbf{\epsilon}}|{\mathbf{x}}_{t}={\mathbf{x}}],$ (13)
and can be approximated by a model ${\mathbf{v}}{\bm{\theta}}({\mathbf{x}}_{t},t)$ trained to minimize the objective:
$\mathcal{L}_{\text{velocity}}=\mathbb{E}_{{\mathbf{x}}_{\ast},\bm{\epsilon},t}\Big{[}||{\mathbf{v}}_{\theta}({\mathbf{x}}_{t},t)-\dot{\alpha}_{t}{\mathbf{x}}_{\ast}-\dot{\sigma}_{t}{\mathbf{\epsilon}}||^{2}\Big{]}.$ (14)
@@
where the score ${\mathbf{s}}({\mathbf{x}}_{t},t)$ is defined as:
${\mathbf{s}}({\mathbf{x}}_{t},t)=-\frac{1}{\sigma_{t}}\mathbb{E}[{\mathbf{\epsilon}}|{\mathbf{x}}_{t}={\mathbf{x}}].$ (16)
The score ${\mathbf{s}}({\mathbf{x}}_{t},t)$ can also be approximated using a model ${\mathbf{s}}_{\bm{\theta}}({\mathbf{x}},t)$, trained with the objective:
$\mathcal{L}_{\text{score}}(\theta)=\mathbb{E}_{{\mathbf{x}}_{\ast},\bm{\epsilon},t}\Big{[}||\sigma_{t}{\mathbf{s}}_{\theta}({\mathbf{x}}_{t},t)+{\mathbf{\epsilon}}||^{2}\Big{]}.$ (17)
@@ Appendix C Hyperparameter and More Implementation Details
More implementation details. We implement our models based on the original DiT and SiT implementation. To speed up training and save GPU memory, we use mixed-precision (fp16) with a gradient clipping and FusedAttention [15] operation for attention computation. We also pre-compute compressed latent vectors from raw pixels via stable diffusion VAE [62] and use these latent vectors. We do not apply any data augmentation, but we find this does not lead to a big difference, as similarly observed in MaskDiT [77] and REPA [74]. We also use stabilityai/sd-vae-ft-ema for encoding images to latent vectors and decoding latent vectors to images. For projection head used for nonlinear transformation, we use two-layer MLP with SiLU activations [19]. The $\lambda$ is set to 0.04 for DiT and 0.2 for SiT initially and followed the rules: $\lambda=\lambda_{\text{init}}*0.1^{\left(\frac{N_{e}-150}{1000}\right)},$ where $N_{e}>150$ is the number of epoch, to change after 150 epochs for ensuring $\mathcal{L}_{\text{gen}}$ and $\mathcal{L}_{\text{sa}}$ at the same scale.
@@
Linear probing. We follow the setup used in REPA [74] and I-DAE [12], and use the code-base of ConvNeXt [50] to conduct the experiment. Specifically, we use AdaptiveAvgPooling and a batch normalization layer to process the latent features output by the model, then train a linear layer for 80 epochs. The batch size is set to 4096 with cosine decay learning rate scheduler, where the initial learning rate is set to 0.001.
@@ Appendix D Evaluation Metric
In this section we explain the main concept of metrics that we used for the evaluation.
  • FID [28] measures the feature distance between the distributions of real and generated images. It uses the Inception-v3 network [66] and computes distance based on an assumption that both feature distributions are multivariate gaussian distributions.
  • sFID [54] proposes to compute FID with intermediate spatial features of the Inception-v3 network to capture the generated images’ spatial distribution.
  • IS [64] also uses the Inception-v3 network but use logit for evaluation of the metric. Specifically, it measures a KL-divergence between the original label distribution and the distribution of logits after the softmax normalization.
  • Precision and recall [41] are based on their classic definitions: the fraction of realistic images and the fraction of training data manifold covered by generated data.
Appendix E Baselines for Comparison
In what follows, we explain the main idea of baseline methods that we used for the evaluation.
  • ADM [17] improves U-Net-based architectures for diffusion models and proposes classifier-guided sampling to balance the quality and diversity tradeoff and improve the performance.
  • VDM++ [37] proposes a simple adaptive noise schedule for diffusion models to improve training efficiency.
  • Simple diffusion [32] proposes a diffusion model for high-resolution image generation by exploring various techniques to simplify a noise schedule and architectures.
  • CDM [30] introduces cascaded diffusion models and trains multiple diffusion models starting from the lowest resolution and applying one or more super-resolution diffusion models for generating high-fidelity images.
  • LDM [63] proposes latent diffusion models by modeling image distribution in a compressed latent space to improve the training efficiency without sacrificing the generation performance.
  • DiT [57] proposes a pure transformer backbone for training diffusion models based on proposing AdaIN-zero modules.
  • SiT [52] extensively analyzes how DiT training can be efficient by moving from discrete diffusion to continuous flow-based modeling.
@@ Appendix E Baselines for Comparison
  • REPA [74] achieves significant improvements in both training efficiency and generation quality by aligning latent feature of diffusion model with that of a large-scale data pre-trained representation model (e.g., DINOv2 [56]).
  • MAETok changes the SD-VAE to MAE-Tok which is trained with auxiliary mask reconstruction loss and align loss with three representation targets (HOG’s [14], DINOv2’s [56], and CLIP’s [60]) and obtains diffusion transformer with better generation performance.
@@ Appendix F Convergence Speed Against REPA
Here we provide quantitative comparison on convergence speed with SiT-XL baseline and SiT-XL + REPA [74]. It is worth noting that REPA leverages a powerful representation foundation model whose training data and resource are far beyond training a diffusion transformer on ImageNet.
As shown in Figure 8, although REPA converges quickly at the beginning of training due to its strong representation prior, the performance saturates after about 200 epochs. On the contrary, in SRA, since the teacher’s ability is constantly enhanced during training and thus better and better guidance is provided, the continuous performance improvement can be achieved. From another perspective, the representation guidance signal in REPA is fixed in every timestep and epoch, while our supervision signal is dynamically changing. This provides a wider variety of learning cues to help the model continuously improve the performance.
@@ Appendix G Detailed Setup of Ablation Study
We now give the detailed experimental setup of Figure 6c in our ablation study. The accuracy rate of the horizontal axis in the figure is obtained by using the linear probing results of the original SiT-XL/2 checkpoint training for 7M iterations, while the vertical axis is the FID evaluation result of training 400K iterations with SRA without classifier-free guidance (CFG). Since we do not introduce any representation component and use the representation supervision signal only in generative training process, we consider this experiment can validate the effectiveness of our approach.
@@ Appendix H Ablation Results of DiT
Figure 9: We also investigate the correlation between generation performance and the representation guidance of DiT + SRA. The similar tight coupling can also be seen.
@@ Appendix I More Discussion on Related Work
Diffusion transformers. Currently, the diffusion model is progressively transitioning from a U-Net-based architecture to a Transformer-based one, owing to the superior scalability of the latter. At the beginning, U-ViT [3] shows transformer-based backbones with skip connections can be an effective backbone for training diffusion models. Then, DiT [57] shows skip connections are not even necessary components, and a pure transformer architecture can be a scalable architecture for training diffusion-based models. Based on DiT, SiT [52] shows the model can be further improved with continuous stochastic interpolants [2]. Moreover, Stable diffusion 3 [20], Lumina-Image 2.0 [59] and FLUX 1 [42] show pure transformers can be scaled up for challenging text-to-image generation, and this characteristic is also verified by Sora [5], CogvideoX [72] and Wan [69] in text-to-video field. Our work fouces on improves the training of DiT (and SiT) architecture based on our proposed self-representation alignment.
@@
Exploring representation capacity of diffusion models. With the success of diffusion models to generate detailed images, many works have attempted to test out whether the discriminative semantic information can be found in diffusion models. GD [53] and DDAE [70] first observe that the intermediate representations of diffusion models have discriminative properties. Driving from this finding, I-DAE [12] deconstructs diffusion models to be a self-supervised Learner. Moreover, Repfusion [71] and DreamTeacher [43] propose knowledge distillation schemes using diffusion models to perform various downstream tasks (e.g., semantic segmentation and object detection). Our work also try to explore representation capacity of diffusion model, but we focus on leveraging the representations in diffusion transformers to enhance their generation capacity.
@@
Applying representation guidance to other generation tasks. In addition to pre-training of the class-conditional diffusion model, applying representation guidance can benefit other generation tasks. For example, lbGen [34] utilizes text feature from CLIP [60] as a low-biased reference to regulate diffusion model for low-biased dataset synthetics. RCG [44] focuses on unconditional generation, which first use the features from self-supervised image encoder to train a representation generator and subsequently used it output as the ‘label’ for image generation by a second generator. Diff-AE [58] uses a learnable encoder for discovering the high-level semantics and a diffusion model for modeling stochastic, these dual-encoding improves realistic of the generated images on attribute manipulation and image interpolation tasks. Different from these works, our study focus on class-conditional diffusion transformer’s pre-training and exploiting representation guidance in itself and it’s own training paradigm. But we also believe our plug-and-play method can be easily applied to other tasks with benefits.

### v3

@@ B.2 Stochastic Interpolants Models (SiT)
${\mathbf{x}}_{t}=\alpha_{t}{\mathbf{x}}_{0}+\sigma_{t}\bm{\epsilon},\quad\alpha_{0}=\sigma_{1}=1,\,\,\alpha_{1}=\sigma_{0}=0,$ (10)
@@
and can be approximated by a model ${\mathbf{v}}_{\bm{\zeta}}({\mathbf{x}}_{t},t)$ trained to minimize the objective:
$\mathcal{L}_{\text{velocity}}=\mathbb{E}_{{\mathbf{x}}_{0},\bm{\epsilon},t}\Big{[}||{\mathbf{v}}_{\bm{\zeta}}({\mathbf{x}}_{t},t)-\dot{\alpha}_{t}{\mathbf{x}}_{0}-\dot{\sigma}_{t}\bm{\epsilon}||^{2}\Big{]}.$ (12)
@@ Appendix C Hyperparameter and More Implementation Details
More implementation details. We implement our models based on the original DiT and SiT implementation. To speed up training and save GPU memory, we use mixed-precision (fp16) with a gradient clipping and FusedAttention [17] operation for attention computation. We also pre-compute compressed latent vectors from raw pixels via stable diffusion VAE [68] and use these latent vectors. We do not apply any data augmentation, but we find this does not lead to a big difference, as similarly observed in MaskDiT [84] and REPA [81]. We also use stabilityai/sd-vae-ft-ema for encoding images to latent vectors and decoding latent vectors to images. For the projection head used for nonlinear transformation, we use two-layer MLP with SiLU activations [21]. The $\lambda$ is set to 0.04 for DiT and 0.2 for SiT initially and followed the rules: $\lambda=\lambda_{\text{init}}*0.1^{\left(\frac{N_{e}-150}{1000}\right)},$ where $N_{e}>150$ is the number of epoch, to change after 150 epochs for ensuring $\mathcal{L}_{\text{gen}}$ and $\mathcal{L}_{\text{sa}}$ at the same scale.
@@
Linear probing. We follow the setup used in REPA [81] and I-DAE [13], and use the code-base of ConvNeXt [53] to conduct the experiment. Specifically, we use AdaptiveAvgPooling and a batch normalization layer to process the output of the latent feature by the model, then train a linear layer for 80 epochs. The batch size is set to 4096 with a cosine decay learning rate scheduler, where the initial learning rate is set to 0.001.
@@ Appendix D Evaluation Metric
In this section, we define the key metrics used to assess the performance of our model. Each metric is summarized below:
  • Fréchet Inception Distance (FID) [31]
    Purpose: Measures the similarity between the feature distributions of real and generated images.
    Methodology: Uses the Inception-v3 network [73] to extract features. Assumes both feature distributions are multivariate Gaussian and computes the Fréchet distance between them.
    Interpretation: Lower FID scores indicate better similarity (and thus higher-quality generated images).
  • Spatial Fréchet Inception Distance (sFID) [59]
    Purpose: Extends FID by incorporating spatial information to better capture the structural fidelity of generated images.
    Methodology: Computes FID using intermediate spatial features (rather than global features) from the Inception-v3 network.
  • Inception Score (IS) [70]
    Purpose: Evaluates the quality and diversity of generated images.
    Methodology: Uses the Inception-v3 network to compute class probabilities (logits) for generated images. Measures the KL-divergence between the marginal class distribution of generated images and the conditional class distribution of a single image (after softmax normalization).
    Interpretation: Higher IS values indicate both high image quality (confident predictions) and diversity (uniform marginal distribution).
  • Precision and Recall for Distributions (Precision/Recall) [44]
    Purpose: Evaluates the trade-off between sample quality (precision) and distribution coverage (recall).
    Methodology: Precision: Fraction of generated images deemed realistic by a classifier (relative to real images). Recall: Fraction of the real image manifold covered by generated samples.
Appendix E Methods for Comparison
Next, we explain the key ideas behind the methods used for evaluation and comparison.
  • ADM [19] enhances U-Net-based architectures for diffusion models and introduces classifier-guided sampling, a technique used to balance the quality-diversity tradeoff and improve overall performance.
  • VDM++ [40] improves diffusion model training efficiency through an adaptive noise scheduling mechanism that dynamically adjusts noise levels during optimization.
  • Simple Diffusion [35] targets high-resolution synthesis by simplifying both network architectures and noise schedules through systematic exploration of lightweight design choices.
  • CDM [33] achieves high-fidelity generation via a cascaded pipeline: training low-resolution diffusion models first, then progressively refining details through super-resolution diffusion stages.
  • LDM [69] accelerates training while maintaining generation quality by operating in compressed latent spaces, modeling image distributions at reduced dimensionality compared to pixel space.
  • DiT [62] employs a pure transformer backbone for diffusion models, incorporating AdaIN-zero modules as an important component modification against vanilla vision transformer [20] of its architecture.
  • SiT [56] provides an extensive analysis of how the training efficiency of DiTs can be improved by transitioning from discrete diffusion to continuous flow-based modeling, which can be seen as a flow-based version of DiT.
@@ Appendix E Methods for Comparison
  • REPA [81] achieves significant improvements in both training efficiency and generation quality by aligning the latent feature of the diffusion model with that of a large-scale data pre-trained representation model (e.g., DINOv2 [61]).
  • MAETok changes the SD-VAE to MAE-Tok, which is trained with auxiliary mask reconstruction loss and aligns loss with three representation targets (HOG’s [15], DINOv2’s [61], and CLIP’s [65]) and obtains diffusion transformer with better generation performance.
@@ Appendix F Convergence Speed Against REPA
Here we provide a quantitative comparison of convergence speed with SiT-XL baseline and SiT-XL + REPA [81]. It is worth noting that REPA leverages a powerful representation foundation model whose training data and resources are far beyond training a diffusion transformer on ImageNet.
As shown in Figure 8, although REPA converges quickly at the beginning of training due to its strong representation prior, the performance saturates after about 200 epochs. On the contrary, in SRA, since the teacher’s ability is constantly enhanced during training and thus better and better guidance is provided, continuous performance improvement can be achieved. From another perspective, the representation guidance signal in REPA is fixed in every timestep and epoch, while our supervision signal is dynamically changing. This provides a wider variety of learning cues to help the model continuously improve its performance.
@@ Appendix G Detailed Setup of Ablation Study
We now give the detailed experimental setup of Figure 6c in our ablation study. The accuracy rate of the horizontal axis in the figure is obtained by using the linear probing results of the original SiT-XL/2 checkpoint training for 7M iterations, while the vertical axis is the FID evaluation result of training 400K iterations with SRA without classifier-free guidance (CFG). Since we do not introduce any representation component and use the representation supervision signal only in the generative training process, we consider this experiment to validate the effectiveness of our approach.
@@ Appendix H Ablation Results of DiT
Figure 9: We also investigate the correlation between generation performance and the representation guidance of DiT + SRA. A similar tight coupling can also be seen.
@@ Appendix I More Discussion on Related Work
We now provide a detailed literature review of other related work.
@@
Diffusion transformers. Currently, the diffusion model is progressively transitioning from a U-Net-based architecture to a Transformer-based one, owing to the superior scalability of the latter. At the beginning, U-ViT [3] shows transformer-based backbones with skip connections can be an effective backbone for training diffusion models. Then, DiT [62] shows skip connections are not even necessary components, and a pure transformer architecture can be a scalable architecture for training denoise-based models. Based on DiT, SiT [56] shows the model can be further improved with continuous stochastic interpolants [2]. Moreover, Stable diffusion 3 [22], Lumina-Image 2.0 [64] and FLUX 1 [45] show pure transformers can be scaled up for challenging text-to-image generation, and this characteristic is also verified by Sora [5], CogvideoX [79] and Wan [76] in text-to-video field. Our work focuses on improving the training of DiT (and SiT) architecture based on our proposed self-representation alignment technique.
@@
Exploring representation capacity of diffusion models. With the success of diffusion models to generate detailed images, many works have attempted to test whether discriminative semantic information can be found in diffusion models. GD [58] and DDAE [77] first observe that the intermediate representations of diffusion models have discriminative properties. Driving from this finding, I-DAE [13] deconstructs diffusion models to be a self-supervised Learner. Moreover, Repfusion [78], DiffusionDet [11], and DreamTeacher [46] use diffusion models to perform various downstream tasks (e.g., semantic segmentation and object detection). Our work also tries to explore the representation capacity of the diffusion model, but we focus on leveraging the representations in diffusion transformers to enhance their generation capacity.
@@
Applying representation guidance to other generation tasks. In addition to pre-training of the class-conditional diffusion model, applying representation guidance can benefit other generation tasks. For example, lbGen [37] utilizes text features from CLIP [65] as a low-biased reference to regulate diffusion model for low-biased dataset synthetics. RCG [47] focuses on unconditional generation, which first use the features from a self-supervised image encoder to train a representation generator and subsequently used it output as the ‘label’ for image generation by a second generator. Diff-AE [63] uses a learnable encoder for discovering the high-level semantics and a diffusion model for modeling stochastic; this dual-encoding improves the realism of the generated images on attribute manipulation and image interpolation tasks. Different from these works, our study focuses on the class-conditional diffusion transformer’s pre-training and exploiting representation guidance in itself and its own training paradigm. But we also believe our plug-and-play method can be easily applied to other tasks with benefits.
Knowledge distillation for diffusion model. Knowledge distillation [27] is also widely used in diffusion models. Its purposes can roughly be divided into two types: improving generation performance and accelerating inference sampling. To achieve the first goal, A more powerful pretrained model is often used to guide the diffusion model during training [16, 81, 24]. For example, TinyFusion [24] combines progressive distillation with architecture-specific optimizations for U-Net backbones, enabling deployment on edge devices. REPA [81] distills the knowledge from a large-scale, pretrained representation foundation model [61, 65] and finds that this distillation can improve both training efficiency and generation quality. Meanwhile, to achieve the second goal, the student and teacher models will be initialized by pre-trained models (here it can be one model or two different models), then the training target is to make the prediction of the student model with fewer steps conform to that of the teacher model with multiple steps [55, 66, 57, 71], thereby achieving the result of accelerated sampling and generation speed. For example, CM [71] constrains the student’s outputs of adjacent points on a sampling path to be consistent with teacher’s. LCM [55] leverages this idea in the latent space. Our study also have some similarities, while we do not need a pretrained model and improve the generation performance of diffusion transformers by applying proposed SRA.
@@
Figure 20: Uncurated samples of laptop (class label: 620).
@@
Figure 23: Uncurated samples of cliff drop-off (class label: 972).